OpenStudy (anonymous):

A positive number N has 3 digits when expressed in base 9. When N is expressed in base 6 the digits are reversed in order. The middle digit is?

OpenStudy (anonymous):

What does base mean here??

OpenStudy (anonymous):

i don't either

OpenStudy (anonymous):

@FoolForMath , Is it related to log ??

OpenStudy (henryblah):

It's like a different counting system. We count in base 10.

OpenStudy (anonymous):

for example base 2 101011100011

OpenStudy (anonymous):

9^2 9^1 9^0 2 0 2 in base 9 is 2(9^2) + 0(9) + 2 (1) =164 in base 10

OpenStudy (anonymous):

@tim538 do you know the answer?

OpenStudy (anonymous):

yeh true but what are the firest digits

OpenStudy (anonymous):

sorry i don't

OpenStudy (anonymous):

121 in base 6 ia 1(6^2) + 2(6) +1(1)=49 in base 10

OpenStudy (anonymous):

all digit of N must be <6

OpenStudy (anonymous):

Okay, say the digits are a, b and c then , 81 a +9b+c = 36 c +6b+a => 80 a+3b-35c =0 5 divides both 80 and 35 so 5 must divides b, hence b =0

OpenStudy (anonymous):

or b = 5

OpenStudy (anonymous):

\[ 0 \le b \lt 6 \]

OpenStudy (anonymous):

\[\huge(abc)_9=(cba)_6\]

OpenStudy (anonymous):

\[81a+9b+c=36a+6b+a\] \[80a+3b=35c\] \[a,b,c<6\]

OpenStudy (anonymous):

2 5 5 in base 9 is 212 un base 10 5 5 2 in base 6 is 212 The answer is 5